OpenStudy (anonymous):

x^3-8 factor the sum

OpenStudy (anonymous):

\[(x^3 - 2^3)\]\[a^3 - b^3 = (a-b)(a^2 + b^2 +ab)\]

OpenStudy (anonymous):

Now you can do it ! : )

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

can you explain. i dont get factoring

OpenStudy (anonymous):

\[\left( x ^{3} - 8 \right)\] \[\left( x-2 \right) \left( x ^{2}+2x+4 \right)\]

OpenStudy (anonymous):

why

OpenStudy (anonymous):

Because if you \[\sqrt[3]{8}\] you get 2 \[2 \times 2 \times 2 = 8\] The formula is: \[\left( a-b \right)\left( a^{2}+ab+b ^{2} \right)\] So you are going to multiply \[x \times x = x ^{2}\] Then you check the sign of your equation. For the second set of parenthesis the first sign is always going to be opposite of your equation and the second sign is always going to be positive Then you multiply a*b Then you multiply b*b \[\left( x ^{3} - 8 \right)\] \[\left( x - 2 \right) \left( x ^{2} + 4x +4 \right)\]